### v1

# Relational Conformal Prediction for Correlated Time Series

Abstract
We address the problem of uncertainty quantification in time series forecasting by exploiting observations at correlated sequences. Relational deep learning methods leveraging graph representations are among the most effective tools for obtaining point estimates from spatiotemporal data and correlated time series. However, the problem of exploiting relational structures to estimate the uncertainty of such predictions has been largely overlooked in the same context. To this end, we propose a novel distribution-free approach based on the conformal prediction framework and quantile regression. Despite the recent applications of conformal prediction to sequential data, existing methods operate independently on each target time series and do not account for relationships among them when constructing the prediction interval. We fill this void by introducing a novel conformal prediction method based on graph deep learning operators. Our method, named CoRel (CoRel), does not require the relational structure (graph) to be known as a prior and can be applied on top of any pre-trained time series predictor. Additionally, CoRel includes an adaptive component to handle non-exchangeable data and changes in the input time series. Our approach provides accurate coverage and archives state-of-the-art uncertainty quantification in relevant benchmarks.
1 Introduction
Many recent advancements in deep learning methods for time series forecasting rely on learning from large collections of (related) time series [1, 2]. In many application domains, such time series are characterized by a rich spatiotemporal dependency structure that can be exploited by introducing inductive biases in the forecasting architecture [3], to steer the learning procedure toward the most plausible models. Accounting for the existing dependencies, usually represented as a graph, allows the resulting models to obtain accurate predictions with a reduced sample complexity [4, 3]. Besides the accuracy of the point estimates, the reliability of the forecasts is a critical aspect of the problem and a key element to enable effective decision-making in many applications [5, 6]. Uncertainty quantification methods [7, 8] can improve reliability by providing confidence intervals on the forecasting error magnitude, allowing for making more informed decisions [9]. This is particularly true for risk-sensitive applications such as healthcare [10] and load forecasting [11]. In this context, inter-series (spatiotemporal) dynamics offer both a challenge and an opportunity. Indeed, while these dependencies can lead to wide prediction intervals (PIs) if overlooked, they may also provide additional knowledge to reduce uncertainty [12].
Existing probabilistic forecasting frameworks often rely on strong distributional assumptions and major modifications of the base point predictor [1, 13]. As such, they cannot be used to quantify uncertainty given a pre-trained forecasting model. In such a setting, conformal prediction (CP) [8, 14] methods are particularly appealing. CP is an uncertainty quantification framework that estimates confidence intervals with marginal coverage guarantees from observed prediction residuals. One of the main assumptions of standard CP methods is that of exchangeability between the data used to estimate the confidence intervals and the test data points, i.e., the assumption that their joint probability distribution is invariant to the ordering of the associated sequence of random variables [15]. Although this assumption does not usually hold when operating on time series [16], several methods have successfully adapted CP to estimate forecast uncertainty [17, 18, 19, 20, 21]. Nevertheless, existing CP approaches operate on each (possibly multivariate) time series independently and cannot account for dependencies among correlated time series.
In this paper, we propose CoRel, a novel CP approach leveraging graph representations and graph deep learning (GDL) for quantifying uncertainty in correlated time series forecasting. In our framework, a spatiotemporal graph neural network (STGNN) [4, 3] is trained on a calibration set to approximate the quantile function of the distribution of prediction residuals. Relationships among time series, assumed to be sparse, are learned end-to-end from the observed residuals owing to a graph structure learning module integrated into the processing. Our approach estimates the error quantile function for each time series at each time step, by conditioning the shared uncertainty quantification model on past observations at neighboring nodes (as defined by the learned graph structure). Finally, an adaptive component is added to handle potential non-stationarities by relying on a small set of parameters specific to each time series. Our approach can be applied to the residuals generated by any point forecasting model, even those that completely disregard potential relationships among the input time series.
Our main novel contributions can be summarized as follows.
  • The first application of GDL to CP for time series;
  • A novel, sound, and effective CP method able to quantify uncertainty from observations across a collection of correlated time series;
  • A family of graph-based architectures to estimate uncertainty that share most of the learnable parameters among the processed time series, while including node-level parameters that dynamically adapt to changes in each target sequence.
Empirical results show that CoRel achieves state-of-the-art performance compared to existing CP approaches for time series in several datasets and under different scenarios.
2 Preliminaries
2.1 Problem formulation
Consider a collection of $N$ sparsely correlated time series. Denote by ${\bm{x}}^{i}_{t}\in{\mathbb{R}}$ the scalar target variable associated with the $i$-th time series at time step $t$; ${\bm{X}}_{t}\in{\mathbb{R}}^{N\times 1}$ indicates the $N$ stacked target variables w.r.t. the entire time series collection. ${\bm{X}}_{t:t+T}$ indicates the sequence within time interval $[t,t+T)$; conversely, with the shorthand ${\bm{X}}_{<t}$ refers to observations up to time step $t$ (excluded). Time series are assumed to be homogenous, i.e., all the variables (observables) describe the same physical quantity (e.g., temperature or energy consumption). Analogously, ${\bm{U}}_{t}\in{\mathbb{R}}^{N\times d_{u}}$ indicates the $d_{u}$-dimensional exogenous covariates associated with each time series. We assume that the $i$-th time series is generated by a stochastic time-invariant process such as
${\bm{x}}_{t}^{i}\sim p\left({\bm{x}}_{t}^{i}|{\bm{X}}_{<t},{\bm{U}}_{<t}\right).$ (1)
Let us hypothesize the existence of a sparse predictive causality à la Granger [22], i.e., we assume that the values of a single time series are related to the values of a (small) subset of other time series in the collection. The extension of the framework to collections of multivariate time series is straightforward, but we focus on the univariate case to simplify the presentations. The problem of dealing with non-stationary processes will be discussed in Sec. 3.4.
Forecasting
We are interested in a model that makes point forecasts by predicting the unknown $H$-steps-ahead ($H\geq 0$) observation ${\bm{X}}_{t+H}$ given a window $W\geq 1$ of past observations ${\bm{X}}_{t-W:t}$ and the associated exogenous variables ${\bm{U}}_{t-W:t}$ as
$\widehat{\bm{X}}_{t+H}={\mathcal{F}}_{\bm{\theta}}({\bm{X}}_{t-W:t},{\bm{U}}_{t-W:t}).$ (2)
${\mathcal{F}}_{\theta}$ denotes a generic parametric model family, i.e., a simple recurrent neural network (RNN) for univariate time series. Given a trained model, our objective is to build a confidence interval around predictions $\widehat{\bm{X}}_{t+H}$. Note that the following easily extends to multi-step predictions $\widehat{\bm{X}}_{t:t+H}$, but we focus on forecasting the single time step $H$ to simplify the presentation and discussion.
Uncertainty quantification
Our objective is to estimate PIs, $C^{\alpha}_{i,t}(\widehat{\bm{X}}_{t+H})$, such that
$P\left({\bm{x}}^{i}_{t}\in C^{\alpha}_{i,t}\left(\widehat{\bm{X}}_{t+H}\right)\right)\geq 1-\alpha,$ (3)
where $\alpha$ is the desired confidence level. If the interval satisfies Eq. 3, we say that the PI achieves marginal coverage $1-\alpha$. Similarly, we say that the PI provides conditional coverage if
$P\left({\bm{x}}^{i}_{t}\in C^{\alpha}_{i,t}\left(\widehat{\bm{X}}_{t+H}\right)\Big{|}{\bm{X}}_{<t},{\bm{U}}_{<t}\right)\geq 1-\alpha.$ (4)
Conditional coverage provides stronger guarantees and it is often harder to achieve [15]. In the following, we will omit the dependence of the interval on the forecasts and simply write $C^{\alpha}_{i,t}$. Among uncertainty quantification methods, we are interested in post-hoc approaches that can build confidence intervals for any given pre-trained point predictor ${\mathcal{F}}_{\theta}$ without requiring any modification of the base forecasting architecture.
2.2 Conformal prediction
As anticipated in Sec. 1, standard CP methods [8, 14] are a class of distribution-free uncertainty quantification techniques that build PIs from empirical quantiles of conformal scores. In the forecasting setting, we consider as conformal scores the prediction residuals,
${\bm{r}}^{i}_{t}={\bm{x}}^{i}_{t}-\hat{\bm{x}}^{i}_{t},$ (5)
and use ${\bm{R}}_{t}$ to denote residuals w.r.t. the entire time series collection. Under appropriate assumptions, CP methods can build valid and informative PIs [14, 16]. SPLIT CONFORMAL PREDICTION (SCP) [8] is arguably the most common approach and exploits scores computed on a calibration set that is disjoint from the training data (i.e., a post-hoc approach).
As mentioned, most standard CP methods rely on the assumption that calibration and test data are exchangeable which allows the procedure to treat them symmetrically and obtain valid (marginal) coverage guarantees [15]. Since this assumption does not hold when dealing with time series data, there have been several recent results extending the CP framework beyond exchangeability [23, 17, 24, 18]. In particular, Barber et al. [16] showed that approximate coverage can be achieved by reweighting the residuals to account for the lack of exchangeability between calibration and test set. Auer et al. [21] learn such a reweighting scheme through an attention-based architecture. Differently, Xu and Xie [19] introduce SCPI, a method based on fitting a quantile random forest [25] on the most recent prediction residuals at each time step. Similar to SCPI, our approach relies on quantile regression to build PIs but differently from existing methods, it exploits observations in arbitrary sets of time series by relying on GDL operators.
2.3 Quantile regression
Quantile regression [26] is an established statistical framework that consists of learning a model of the quantile function (the inverse c.d.f.) of a target distribution from observations. In particular, given ${\bm{y}}\sim p({\bm{y}}|{\bm{x}})$ and observations $(x_{1},y_{1}),\dots,(x_{N},y_{N})$, a standard approach to estimate the $\alpha$-quantile is to train a model by minimizing the so-called pinball loss
$\ell^{\alpha}(\hat{q}^{\alpha}(x),y)=\begin{cases}(1-\alpha)(\hat{q}^{\alpha}(x)-y),\,&\hat{q}^{\alpha}(x)\geq y\\ \alpha(y-\hat{q}^{\alpha}(x)),\,&\hat{q}^{\alpha}(x)<y\end{cases}$ (6)
where $\hat{q}^{\alpha}(x)$ is the estimate of the $\alpha$-quantile w.r.t. $x$.
Quantile networks
Quantile regression has been incorporated in several probabilistic forecasting architectures [1]. The simplest approach consists of using a multi-output network to predict a set of quantiles of interest and interpolate among them to approximate the entire quantile function [27]. More complex approaches rely on, e.g., splines [28]. Conversely, implicit quantile networks (IQNs) [29, 30, 31] approximate the quantile function by being trained to minimize the loss in Eq. 6 given the quantile level $\alpha$ as input and sampling a random $\alpha$ for each sample in a mini-batch.
Figure 1: Overview of CoRel. Past residuals are used as input to a hybrid global-local graph-based quantile network.
2.4 Graph deep learning for time series forecasting
Graph neural networks (GNNs) [32, 33] process graph-structured data by incorporating the graph topology as an inductive bias, e.g., by relying on message-passing layers [34]. STGNNs [4, 3] leverage message-passing layers within sequence modeling architectures to process spatiotemporal data and collections of time series where dependencies are represented as a (possibly dynamic) graph. We consider as reference architectures time-then-space (TTS) models [35, 3] where each time series in the collection is processed independently from the others by a temporal encoder whose output is then fed into a stack of GNN layers. In particular, we adopt the following template architecture:
${\bm{h}}^{i,0}_{t}$ $=\textsc{SeqEnc}\left({\bm{x}}^{i}_{t-W:t},{\bm{u}}^{i}_{t-W:t}\right),$ (7)
${\bm{H}}_{t}^{l+1}$ $=\text{GNN}_{l}({\bm{H}}_{t}^{l},{\bm{A}}),,\quad l=0,\dots,L-1$ (8)
$\hat{\bm{y}}_{t}$ $=\textsc{Readout}\left({\bm{h}}_{t}^{L}\right),$ (9)
where ${\bm{A}}\in{\mathbb{R}}^{N\times N}$ is the graph adjacency matrix and $\hat{\bm{y}}_{t}$ a generic node-level prediction associated with the problem at hand. $\textsc{SeqEnc}({}\cdot{})$ and $\text{GNN}_{l}({}\cdot{})$ denote, respectively, any sequence modeling architecture, e.g., an RNN, and any GNN layer, e.g., based on message-passing. Representations can then be mapped into predictions $\widehat{\bm{Y}}_{t}$ by using any readout block, e.g., a multilayer perceptron (MLP). STGNNs have been used as forecasting architecture (${\bm{Y}}_{t}={\bm{X}}_{t+H}$) with great success. In the following, we will exploit this framework as a backbone for estimating the residual quantile distribution. We refer to Sec. 4 and Jin et al. [4] for more discussion on the application of STGNNs in the context of time series analysis.
3 Conformal relational prediction
Our objective is to build PIs by exploiting relational dependencies across the residuals of the target time series. We model the dependencies as edges of a graph and learn them under the assumption that the relational structure is sparse, which reduces the computational costs and act as an inductive bias on the structure learning architecture. By relying on such representation, we can leverage GDL methods for time series to process the data. In particular, we train a STGNN on the residuals of the calibration set to predict the quantiles of the error distribution. Conditioning the prediction on the recent history of related time series allows for taking the dependency structure of the data into account when estimating uncertainty: a key aspect in applying conformal prediction to non-exchangeable data [16]. Compared to existing methods [19, 21] that only capture temporal dependencies, our approach allows for modeling spatiotemporal dependencies among different time series. Sec. 3.1 presents the details of the proposed conformal inference procedure by assuming that the relational structure at each time step is defined by an adjacency matrix ${\bm{A}}\in{\mathbb{R}}^{N\times N}$ (Sec. 3.1). We then show how to learn the graph structure directly from data and make the model adaptive in Sec. 3.2 and Sec. 3.4, respectively. Finally, we discuss the theoretical properties of the approach in Sec. 3.3. Fig. 1 shows an overview of the architecture.
3.1 Relational Quantile Predictor
Consider a standard SCP setup, where the training data are split into training and calibration sets. For the moment, we disregard possible nonstationarities in the data considering the problem setup introduced in Sec. 2.1 and encode spatial dependencies in the adjacency matrix ${\bm{A}}\in{\mathbb{R}}^{N\times N}$. While the training set is used to fit the point predictor ${\mathcal{F}}_{\theta}$, we use the prediction residuals in the calibration set (${\mathcal{R}}^{cal}$) to learn the quantile function of the error distribution at each time step.
We implement the quantile regressor as a hybrid global-local STGNN, which mixes global (shared) parameters with local, target-specific components [36]. Sharing most learnable parameters across all time series reduces sample complexity, while local parameters allow for processing each series differently. Specifically, we keep all processing blocks shared and associate a learnable node embedding ${\bm{v}}^{i}\in{\mathbb{R}}^{d_{v}}$ with each time series [37]. More specifically, our model is a quantile network (see Sec. 2.3) composed of the following processing layers:
${\bm{h}}^{i,0}_{t}$ $=\textsc{Enc}\left({\bm{r}}^{i}_{t-1},{\bm{v}}^{i}\right),$ (10)
${\bm{Z}}_{t}$ $=\textsc{STGNN}\Big{(}{\bm{H}}^{0}_{\leq t},{\bm{A}}\Big{)},$ (11)
$\hat{\bm{q}}^{\alpha,i}_{t+H}$ $=\textsc{QDec}\left(\alpha,{\bm{z}}_{t}^{i},{\bm{v}}^{i}\right),$ (12)
where ${\bm{r}}^{i}_{t-1}$ are prediction residuals (Eq. 5) and $\hat{\bm{q}}^{\alpha,i}_{t+H}$ is the predicted $\alpha$-quantile at time step $t+H$ for the $i$-th time series. $\textsc{Enc}({}\cdot{})$ denotes any encoding layer,e.g., a linear transformation or an MLP. For the STGNN block, several designs are possible (e.g., see Jin et al. [4]); the one we follow is the template in Sec. 2.4. $\textsc{QDec}({}\cdot{})$ is a readout mapping the representations at each node to the prediction of the quantile of specified level $\alpha$. We refer to the family of quantile networks defined in Eq. 1012 as relational quantile predictors (RelQPs) and use the notation
$\widehat{\bm{Q}}^{\alpha}_{t}={\mathcal{Q}}_{\psi}\left(\alpha,{\bm{V}};{\bm{R}}_{t-W:t},{\bm{A}}\right),$ (13)
where ${\mathcal{Q}}_{\psi}$ indicates the shared (global) part of the network and $\widehat{\bm{Q}}^{\alpha}_{t}\in{\mathbb{R}}^{N}$ denotes the predicted $\alpha$-quantiles at time step $t$ w.r.t. the full time series collection. Note that the framework can easily accommodate further inputs at the encoding block (e.g., we can condition the regression on ${\bm{X}}_{<t}$ and ${\bm{U}}_{<t}$). The model is trained by, at each time step in the calibration set, the pinball loss (Eq. 6) w.r.t. the full-time series collection. Through the message-passing layers, the residuals of each time series contribute to estimating the quantiles of the error distribution at neighboring nodes. In practice, we restrict the input of the regressor to the most recent observations rather than considering the full sequence (the window length here can also be different from the one used by the point predictor).
Building the confidence intervals
Given the trained quantile network ${\mathcal{Q}}_{\psi}$, we build the PIs for each target (test) time step as
$\widehat{C}^{\alpha}_{t}=\left[\widehat{\bm{X}}_{t+H}+\widehat{{\bm{Q}}}^{\alpha/2}_{t},\widehat{\bm{X}}_{t+H}+\widehat{{\bm{Q}}}^{1-\alpha/2}_{t}\right],$ (14)
or
$\hat{\beta}$ $=\operatorname*{arg\,min}_{\beta}\Big{|}\widehat{{\bm{Q}}}^{1-\alpha/2+\beta}_{t}-\widehat{{\bm{Q}}}^{\alpha/2+\beta}_{t}\big{|}$ (15)
$\widehat{C}^{\alpha}_{t}$ $=\left[\widehat{\bm{X}}_{t+H}+\widehat{{\bm{Q}}}^{\alpha/2+\hat{\beta}}_{t},\widehat{\bm{X}}_{t+H}+\widehat{{\bm{Q}}}^{1-\alpha/2+\hat{\beta}}_{t}\right],$ (16)
where $\widehat{C}^{\alpha}_{t}$ indicates the estimated PI. While both Eq. 14 and Eq. 16 correspond to the same confidence level, the PI in Eq. 16 can be narrower, at the expense of the additional computation needed to obtain $\hat{\beta}$ [18]. In practice, one can choose between the two approaches given computational constraints and the difference in performance observed on a validation set. Note that $\widehat{C}^{\alpha}_{i,t}$ can provide only approximate coverage, as it is subject to approximation errors of the true quantile function. Sec. 3.3 will discuss this aspect in detail. A potential drawback of CoRel is that residuals cannot be assumed exchangeable in most practical scenarios. The error distribution can be non-stationary, making it difficult to obtain any coverage guarantee. To mitigate the problem, Sec. 3.4 discusses an efficient and scalable approach to make the framework adaptive by updating local components of the architecture over time. Finally, it is worth noting that the use of relational components in CoRel relies on the actual presence of the associated dependencies in the data. In practical applications, the presence of spatiotemporal correlations in the residuals can be verified through ad-hoc statistical tests  [12, 38], whose outcome can support the adoption of CoRel.
3.2 Learning the Relational Structure
Assuming the dependency structure across time series to be unknown, we integrate a graph learning module into the architecture to derive the operational graph topology directly from the residuals. To do so, we adopt a probabilistic structure learning framework [39, 40, 41]. In particular, we associate each edge with a score $\phi^{ij}$ and learn a distribution over $K$-NN graphs parametrized by the matrix ${\bm{\Phi}}\in{\mathbb{R}}^{N\times N}$ [40, 42]. Notably, we consider graphs obtained by sampling, for each $i$-th node, $K$ elements without replacement from the categorical distribution
${\bm{\Phi}}$ $={\mathcal{E}}_{\bm{\xi}}\left({\bm{R}}_{<t},{\bm{V}},\dots\right)$ (17)
$M_{i}$ $=\text{Categorical}\left(\frac{\exp\{\phi^{ik}\}}{\sum_{j=1}^{N}\exp\{\phi^{ij}\}};k\in\{1,\dots,N\}\right),$ (18)
where ${\mathcal{E}}_{\bm{\xi}}({}\cdot{})$ is a generic trainable encoder with parameters $\bm{\xi}$. In practice, sampling can be done efficiently by exploiting the GumbelTopK trick [43] and scores ${\bm{\Phi}}$ can be parametrized directly as ${\bm{\Phi}}=\bm{\xi}$.
End-to-end learning
To propagate gradients through the sampling, we rely on the continuous relations introduced by  Xie and Ermon [44] paired with a straight-through gradient estimator [45] to obtain discrete samples. Optionally, we sparsify the gradients by backpropagating only through a random subset of the zero entries of ${\bm{A}}$ (more details are provided in App. D).
3.3 Theoretical analysis and further discussion
We start the discussion by providing an intuitive bound on the approximate coverage provided by CoRel.
Proposition 3.1.
Let $P^{c}_{t+H}({\bm{X}}_{t+H})=p_{t+H}({\bm{X}}_{t+H}\mid{\bm{X}}_{<t},{\bm{U}}_{<t})$ and $P^{c}_{\psi}({\bm{X}}_{t+H})=p_{\psi}({\bm{X}}_{t+H}\mid{\bm{X}}_{<t},{\bm{U}}_{<t})$ be the true conditional data-generating distribution at the test point $t+H$ and the probability distribution associated with the learned quantile function ${\mathcal{Q}}_{\psi}$, respectively. Then
$P^{c}_{t+H}\left({\bm{X}}_{t+H}\in\hat{C}^{\alpha}_{t}(\widehat{\bm{X}}_{t+h})\right)\geq 1-\alpha-{TV}\left(P^{c}_{\psi},P^{c}_{t+H}\right)$
where ${TV}({}\cdot{})$ denotes the total variation function.
The proof relies on the properties of the total variation of probability measures and can be found in App. A. Here, differently from the problem settings introduced in Sec. 2.1, we do not assume the process to be time-invariant. 3.1 links the conditional coverage gap to the approximation error in estimating the quantile function of the residuals. By making assumptions on the expressivity of the quantile regressor in Eq. 13 and on the stationarity process (e.g., by assuming a strongly mixing process), we can expect the total variation between the learned and true distribution to shrink asymptotically as the size of the calibration set increases. Moreover, in this case, monitoring the coverage gap on a validation set offers an estimate of the actual miscoverage on test data. Conversely, if we expect the process to be non-stationary, ${\mathcal{Q}}_{\psi}$ has to be updated over time to keep the coverage gap contained. Within this context, the next section discusses a simple and sample-efficient approach to make CoRel adaptive. Finally, the bound provided in 3.1 shares similarities with the one in [16], which bounds the miscoverage gap for CP from weighted empirical quantiles. 3.1 can be seen as an analogous result that holds when estimates obtained from empirical quantiles are replaced with a parametric function approximator.
3.4 Adaptativity
The RelQP model introduced in Sec. 3.1 can yield arbitrarily large coverage gaps in the presence of distribution shifts from the calibration set, where the model is trained. Adopting a re-training approach such as in Xu and Xie [19] would be impractical due to the higher sample complexity entailed by the deep learning approach that we adopt. Therefore, to mitigate this issue while keeping the computational complexity under control, we update only the local components of the model over time, i.e., the learnable node embeddings ${\bm{V}}$ [37]. This allows for keeping most of the parameters of the model fixed and fine-tuning only a small number of weights for each node. Empirically, we show that this procedure can effectively improve the quality of the uncertainty estimates.
4 Related Work
The problem of quantifying forecast uncertainty is central in fundamental and applied research in time series forecasting [9, 6]. Among deep learning approaches [1], many generative architectures have been proposed as means to obtain probabilistic forecasts [13, 46, 47, 48]. Most related to our approach are those methods that exploit quantile regression to produce probabilistic forecasts [27, 28, 49, 31]. Similarly to CoRel, these quantile regression techniques do not usually require strong assumptions on the underlying data distribution. Regarding probabilistic graph-based forecasting architecture, the existing literature is limited [4, 3]. Wu et al. [50] investigate the combination of STGNNs with standard uncertainty quantification techniques for deep learning. Pal et al. [51] use an STGNN to implement a state-space model and quantify uncertainty within a Bayesian framework. Wen et al. [52] propose a probability predictor based on combining STGNNs with a diffusion model [53]. Zambon et al. [54] introduce a framework for designing probabilistic graph state-space models that can process collections of time series. However, all these methods cannot operate on top of an existing pre-trained model and require training an ad-hoc forecasting model. Conversely, CoRel is trained, within a CP framework, on predicting the quantiles of the error distribution of an existing model, rather than on forecasting the target variable.
Conformal prediction
Related work on CP for time series has been already discussed in Sec. 2.2 and Sec. 3. Related to our method, Mao et al. [55] propose a CP approach for (static) spatially correlated data. Jiang et al. [56] propose to quantify the uncertainty in predicting power outages by fitting a quantile random forest [25] on time series from neighboring geographical units. CoRel can be framed among the CP methods that learn a model of conformal scores distribution [19, 57]. Differently from existing methods that operate on each time series separately, the estimates are conditioned on errors at both the target time series as well as at neighboring nodes. To the best of our knowledge, no previous CP method has been designed to specifically operate on collections of correlated time series and exploit graph deep learning operators. CP methods for multivariate time series do exist [58, 59], but operate on a single multidimensional time series. Moreover, although global-local models are popular among forecasting architectures [36, 1], CoRel is the first CP architecture of this kind. Finally, CP methods have also been applied to static graphs and used to quantify the uncertainty of GNNs, both in inductive [60] and transductive settings [61]. More recently, Davis et al. [62] proposed a CP method for node classification with GNNs in dynamic networks. These methods often assume node/edge exchangeability [60, 61] or are limited to node classification tasks [63].
Table 1: Performance comparison for $\alpha=0.1$. $\Delta$Cov values are color-coded to highlight different levels of undercoverage: green ($<$2%), yellow (2-3%), orange (3-4%), red ($>$4%). The lowest Winkler score for each scenario is shown in bold.
Metric SCP NexCP SeqCP SCPI HopCPT CoRNN CoRel
METR-LA RNN $\Delta$Cov -0.65 -1.07 -8.10 -0.58$\pm$0.01 -2.04$\pm$0.30 -1.71$\pm$0.60 -1.24$\pm$0.73
PI-Width 20.60 24.12 18.59 19.81$\pm$0.01 17.18$\pm$0.11 17.51$\pm$0.42 16.10$\pm$0.70
Winkler 40.22 39.19 48.29 37.27$\pm$0.01 28.41$\pm$0.20 33.90$\pm$0.21 25.55$\pm$0.44
Transf $\Delta$Cov -0.72 -1.22 -12.23 -0.59$\pm$0.01 -2.69$\pm$0.65 -1.24$\pm$0.39 -1.02$\pm$0.59
PI-Width 20.51 24.21 16.99 19.90$\pm$0.00 16.84$\pm$0.21 17.78$\pm$0.36 15.72$\pm$0.15
Winkler 40.29 39.23 49.07 37.66$\pm$0.01 28.58$\pm$0.36 33.99$\pm$0.21 25.27$\pm$0.15
STGNN $\Delta$Cov -1.61 -1.24 -11.84 -1.56$\pm$0.01 -2.31$\pm$0.29 -0.62$\pm$0.40 -1.54$\pm$1.01
PI-Width 17.68 22.20 13.68 16.59$\pm$0.01 16.07$\pm$0.11 15.94$\pm$0.35 15.40$\pm$0.52
Winkler 36.78 34.34 39.73 34.37$\pm$0.02 26.18$\pm$0.13 30.11$\pm$0.08 26.86$\pm$0.26
CER-E RNN $\Delta$Cov -3.42 0.12 -3.52 -3.41$\pm$0.01 -4.50$\pm$0.59 -1.31$\pm$0.44 -3.89$\pm$0.15
PI-Width 2.61 3.26 2.77 2.40$\pm$0.00 2.07$\pm$0.05 2.11$\pm$0.03 1.88$\pm$0.01
Winkler 5.71 5.46 5.80 5.31$\pm$0.00 4.12$\pm$0.06 4.13$\pm$0.03 3.82$\pm$0.03
Transf $\Delta$Cov -3.30 0.11 -3.60 -3.23$\pm$0.02 -4.51$\pm$0.53 -1.23$\pm$0.90 -3.34$\pm$0.34
PI-Width 2.55 3.18 2.70 2.37$\pm$0.00 2.04$\pm$0.07 2.08$\pm$0.04 1.90$\pm$0.02
Winkler 5.67 5.40 5.76 5.29$\pm$0.00 4.10$\pm$0.09 4.26$\pm$0.04 3.80$\pm$0.04
STGNN $\Delta$Cov -4.41 0.04 -3.63 -4.32$\pm$0.00 -5.28$\pm$0.17 -2.08$\pm$0.12 -5.03$\pm$0.52
PI-Width 2.29 2.98 2.42 2.10$\pm$0.00 1.84$\pm$0.01 1.90$\pm$0.03 1.80$\pm$0.04
Winkler 5.11 4.89 5.00 4.81$\pm$0.00 3.68$\pm$0.02 3.84$\pm$0.04 3.82$\pm$0.05
AQI RNN $\Delta$Cov 4.77 0.04 -2.34 3.80$\pm$0.02 3.97$\pm$1.07 0.47$\pm$1.60 -2.46$\pm$1.64
PI-Width 131.17 86.95 78.87 120.77$\pm$0.05 106.20$\pm$6.45 78.00$\pm$4.16 72.30$\pm$3.56
Winkler 159.23 132.59 136.64 152.17$\pm$0.05 133.87$\pm$2.39 111.21$\pm$2.11 109.79$\pm$1.53
Transf $\Delta$Cov 4.77 0.03 -2.23 3.85$\pm$0.01 4.21$\pm$0.43 -0.82$\pm$1.14 -1.04$\pm$1.07
PI-Width 131.91 87.45 80.15 122.81$\pm$0.03 108.38$\pm$2.70 75.10$\pm$1.67 74.44$\pm$1.47
Winkler 160.66 134.27 138.95 154.63$\pm$0.02 135.54$\pm$1.15 111.11$\pm$0.04 110.15$\pm$1.06
STGNN $\Delta$Cov 5.19 -0.01 -2.69 4.53$\pm$0.01 4.52$\pm$0.21 -0.74$\pm$0.61 -2.00$\pm$1.30
PI-Width 135.89 92.93 81.20 124.27$\pm$0.04 112.72$\pm$1.53 75.97$\pm$0.26 73.51$\pm$2.40
Winkler 163.12 139.80 141.37 154.26$\pm$0.04 138.96$\pm$0.58 112.90$\pm$0.29 112.62$\pm$0.91
5 Experiments
We validate CoRel across three experimental settings. In the first one (Sec. 5.1), we compare it against state-of-the-art CP methods operating on the residuals produced by different forecasting models. Then, we analyze CoRel in a controlled environment (synthetic dataset). Finally, we assess the effectiveness of the procedure described in Sec. 3.4 in adaptively improving the PIs. We implement CoRel as an RNN followed by two message-passing layers. To approximate the quantile function, we train the model by minimizing the pinball loss over a discrete set of quantiles, similarly to Wen et al. [27]. PIs are constructed as in Eq. 14; App. G shows results for the alternative construction in Eq. 16. To learn the graph, we directly parametrize the score matrix ${\bm{\Phi}}$ by associating a learnable parameter with each of its entries. We use as metrics: 1) the difference between the specified confidence level $1-\alpha$ and the observed coverage on the test set ($\Delta$Cov); 2) the width of the PI (PI-Width); 3) the Winkler score [64], which is computed as the width of the PI plus penalty for each observation outside of the predicted interval proportional to the actual error (Winkler). More details and results are provided in the appendix.
5.1 Time series forecasting benchmarks
We consider the following datasets, each coming from a different application domain: METR-LA from the traffic forecasting literature [65]; a collection of air quality measurements from different Chinese cities (AQI[66]; a collection of energy consumption profiles acquired from smart meters within the CER smart metering project (CER-E[67, 68]. We follow the preprocessing steps of previous works [65, 69, 37] and adopt $40\%/40\%/20\%$ splits for training, calibration, and testing, respectively. For each dataset, we first train $3$ different baseline models: a simple RNN with gated recurrent unit (GRU) cells [70], a decoder-only Transformer [71], a simple TTS STGNN by following the template in Sec. 2.4. The latter uses a pre-defined graph that models the dependencies across the time series. After training, we evaluate each baseline on the calibration set and save the associated residuals, which are then used as input to the different CP methods. More details on the datasets and base models are provided in App. C.
Baselines
We compared CoRel against the following methods: 1) SCP, the standard split CP; 2) SeqCP, where, analogously to Xu and Xie [18], we compute empirical quantiles using only the most recent $K$ residuals at each time step; 3) NexCP [16], which computes empirical quantiles by assigning exponentially decaying weights to past residuals; 4) SCPI [19], which estimates the residuals’ quantile function from the last few steps of each time series with a random forest; 5) HopCPT which reweights past residuals by learning attention scores with a Modern Hopfield Network [72]. We include in our comparison a model called CoRNN, where we use the same architecture as CoRel but remove the message-passing layers. Except for HopCPT which uses a custom procedure [21], model selection is performed on a validation set by optimizing the Winkler score.
Table 2: Adaptation results for CER-E for a reference CoRel model. ($\alpha=0.1$).
RNN Transformer STGNN
$\Delta$Cov PI-Width Winkler $\Delta$Cov PI-Width Winkler $\Delta$Cov PI-Width Winkler
W/o adapt. -3.46$\pm$ 0.16 1.91$\pm$ 0.01 3.83$\pm$ 0.05 -3.07$\pm$ 0.31 1.95$\pm$ 0.03 3.82$\pm$ 0.02 -5.15$\pm$ 0.44 1.79$\pm$ 0.03 3.79$\pm$ 0.04
W/ adapt. -2.65$\pm$ 0.12 1.93$\pm$ 0.01 3.61$\pm$ 0.05 -2.39$\pm$ 0.37 1.96$\pm$ 0.01 3.64$\pm$ 0.02 -3.13$\pm$ 0.28 1.85$\pm$ 0.01 3.59$\pm$ 0.03
Results
Tab. 1 reports the results across the datasets and the base prediction models. The first observation is that CoRel outperforms the competitors in terms of Winkler score in almost all cases. We observed a few exceptions only when the baseline is itself an STGNN, as it is already expected to take care of modeling spatiotemporal dependencies. However, note that the STGNN base model has access to a pre-defined graph, which is not always available in practical applications. In terms of coverage, CoRel achieves good results, with the exception of some cases in the CER-E dataset. However, we note that our model selection prioritized the Winkler score which emphasizes the width of the prediction bands besides the coverage. CoRNN, the simplified version of our approach, obtains good overall performance and is competitive against the state-of-the-art but, despite providing good coverage, it is outperformed by CoRel in most scenarios in terms of Winkler score. Among the competitors, HopCPT provides competitive results in METR-LA and CER-E but with a larger coverage gap. Except for SeqCP, the other baselines obtain good coverage in most settings at the expense of drastically wider PIs.
Table 3: Performance on GPVAR.
GPVAR-RNN GPVAR-STGNN
Models $\Delta$Cov Winkler $\Delta$Cov Winkler
SCP -0.023 2.14 -0.019 1.66
CoRNN -0.1$\pm$0.2 2.04$\pm$0.00 0.0$\pm$0.0 1.66$\pm$0.00
CoRel 0.0$\pm$0.1 1.67$\pm$0.00 0.1$\pm$0.1 1.66$\pm$0.00
w/ true ${\bm{A}}$0.1$\pm$0.1 1.66$\pm$0.00 0.0$\pm$0.2 1.66$\pm$0.00
5.2 Controlled environment
We evaluated the behavior of CoRel in a controlled environment by simulating a diffusion process on a graph. In particular, the experiment relied on the GPVAR benchmark introduced by Zambon and Alippi [12], with a setup analogous to [37]. Data were generated recursively from the auto-regressive polynomial graph filter:
${\bm{H}}_{t}$ $=\sum_{l=1}^{L}\sum_{q=1}^{Q}\Theta_{q,l}{\bm{A}}^{l-1}{\bm{X}}_{t-q},$
${\bm{X}}_{t+1}$ $={\bm{a}}\odot\text{tanh}\left({\bm{H}}_{t}\right)+{\bm{b}}\odot\text{tanh}\left({\bm{X}}_{t-1}\right)+\eta_{t},$ (19)
where parameters $\bm{\Theta}\in{\mathbb{R}}^{Q\times L}$, ${\bm{a}}\in{\mathbb{R}}^{N}$, ${\bm{b}}\in{\mathbb{R}}^{N}$ are kept fixed across nodes and $\eta_{t}\sim\mathcal{N}(\bm{0},\sigma^{2}{\mathbb{I}})$. We ran the simulation on a graph with $60$ nodes and with a topology analogous to previous works [37]; more details are provided in App. C. We use both a RNN and STGNN as base point predictors. As shown by Zambon and Alippi [12], STGNNs can obtain a forecasting accuracy near the theoretical optimum in this dataset, which results in i.i.d. residuals. As such, we would expect standard SCP to be sufficient in this scenario. The objective of this experiment is to show that CoRel is effectively able to capture and leverage existing spatiotemporal dependencies.
Results
We compare CoRel against standard SCP and the CoRNN variant. Moreover, we also compare performance against CoRel with access to the true graph used to generate the data. Results are shown in Tab. 3. When using an RNN as a point predictor (GPVAR-RNN) CoRel significantly outperforms both standard SCP and CoRNN. Furthermore, CoRel achieves results that closely match those obtained with direct access to the ground truth graph, showing the effectiveness of the proposed architecture in capturing latent relational dependencies. Finally, results obtained by using an STGNN as baseline (GPVAR-STGNN) show (as expected) that the standard SCP is sufficient when the point-predictor captures all the relevant dependencies.
5.3 Adaptability
In this experiment, we evaluate how effectively the adaptation technique proposed in Sec. 3.4 provides accurate PIs over time. We focused on the CER-E dataset where the calibration set does not cover a full year, which likely introduces a shift at test time given the seasonality of energy consumption. This behavior is also confirmed by the results in Sec. 5.1.
We used CoRel with a fixed hyperparameter configuration across scenarios and trained on the calibration set. We then split the test set in $K=6$ folds and then iteratively evaluated and fine-tuned the model on each fold. Results, reported in Tab. 2, show that the proposed adaptation schemes improve performance by reducing the coverage gap and improving the accuracy of the PI.
6 Conclusion
In this paper, we introduced CoRel (CoRel), a novel CP method for correlated time series. CoRel exploits graph-based neural operators to implement an uncertainty quantification architecture that can operate on top of any pre-trained point predictor. Furthermore, our approach does not require the relational structure to be known in advance. Results show that the proposed method compares favorably against the state-of-the-art in several relevant scenarios.
Future works
We believe that CoRel constitutes an important step toward effective spatiotemporal CP methods. There are several directions for future works to explore, such as applying the framework to heterogenous time series. Future work should also further focus on mitigating issues coming from learning in non-stationary environments.
Appendix
Appendix A Proof of 3.1
As stated in 3.1, let $P^{c}_{t+H}$ and $P^{c}_{\psi}$ indicate the probability distributions associated with the true data distribution at $t+H$ and with the learned quantile function ${\mathcal{Q}}_{\psi}$, respectively. It follows from the definition of total variation distance $\,{TV}(P,Q):=\sup_{B}|P(B)-Q(B)|,$ that for any event $B$
$|P^{c}_{t+H}(B)-P^{c}_{\psi}(B)|\,\leq\,{{TV}}(P^{c}_{\psi},P^{c}_{t+H})\quad\Longrightarrow$ (20)
$P^{c}_{t+H}(B)\,\geq\,P^{c}_{\psi}(B)\,-\,{{TV}}(P^{c}_{\psi},P^{c}_{t+H}).$ (21)
For $\hat{C}^{\alpha}_{t}$ in both Eq. 14 and Eq. 16 we have by construction that
$G^{c}_{\psi}({\bm{X}}_{t+H}\in\hat{C}^{\alpha}_{t})=1-\alpha.$ (22)
Then, putting together Eq. 21 and Eq. 22 we obtain
$P^{c}_{t+H}\left({\bm{X}}_{t+H}\in\hat{C}^{\alpha}_{t}(\widehat{\bm{X}}_{t+h})\right)\geq 1-\alpha-{TV}\left(P^{c}_{\psi},P^{c}_{t+H}\right).$
Hence, the proof is complete.
Appendix B Hardware and software platforms
Benchmarks have been developed with Python [73] and the following open-source libraries:
  • Numpy [74];
  • PyTorch [75];
  • PyTorch Lightning [76];
  • PyTorch Geometric [77];
  • Torch Spatiotemporal [78].
Experiments were conducted on a server equipped with AMD EPYC 7513 CPUs and NVIDIA RTX A5000 GPUs. The code for reproducing the computational experiments will be open-sourced upon publication. For the HopCPT, SCPI, NexCP, and SeqCP baselines we use the open-source implementation made available by Auer et al. [21] [^1] https://github.com/ml-jku/HopCPT.
Appendix C Datasets
We considered several datasets from different application domains, real-world scenarios and a simulated controlled environment. For the real world datasets we followed the pre-processing steps adopted by Cini et al. [37].
Air Quality Monitoring
The AQI [66] dataset consists of 8,760 hourly measurements of pollutant PM2.5 from 437 monitoring stations in China. We use window of $W=24$ time steps and predict the $3$-time-step-ahead observations. For the STGNN base model, we build an adjacency matrix using a thresholded Gaussian kernel computed from pairwise geographic distances [79].
Traffic Forecasting
We considered the METR-LA [65] traffic forecasting dataset, consisting of 34,272 timesteps of measurements from 207 loop detectors sampled at 5-minute intervals in the Los Angeles County Highways. We use window of $12$ time steps and predict the $12$-time-step-ahead observations. For the STGNN base model, we followed previous works [69] and build an adjacency matrix using a thresholded Gaussian kernel applied to geographic distances.
Electric Load Forecasting
We selected the CER-E dataset [67, 68], comprising 25,728 timesteps of energy consumption readings aggregated at 30-minute intervals from 485 smart meters monitoring small and medium-sized enterprises. We use window of $48$ time steps and predict the $5$-time-step-ahead observations. For the STGNN base model, we built the adjacency matrix by extracting a 10-nearest neighbor graph from week-wise correntropy similarities between time series, following previous work [68].
GPVAR
For the GPVAR dataset, we generate synthetic data with 40,000 timesteps over an undirected network of 60 nodes connected in a community graph structure by following the system model in Sec. 5.2 [12]. The parameters of the spatiotemporal process are set as
$\Theta=\begin{bmatrix}2.5&-2.0&-0.5\\ 1.0&3.0&0.0\end{bmatrix},\quad a=b=0.5.$ (23)
We used an input window of $5$ time steps to predict the next observation. For the STGNN base model, we used the same community graph structure as the adjacency matrix.
Base models
We trained three base models (point predictors) for each dataset: a RNN with GRU cells (1 layer with hidden size 32), a decoder-only Transformer (hidden size 32, feed-forward size 64, 2 attention heads, 3 layers, dropout 0.1), and a STGNN following the template in Sec. 2.4 (hidden size 32, node embedding size 16, 1 layer GRU, 2 message-passing layers). All models were trained by minimizing the MAE loss using the Adam optimizer for 200 epochs with batch size 32, using 40% of the data for training, 40% for calibration, and 20% for testing. Input features were standardized across the spatial and temporal dimensions. For the STGNN, we used the pre-defined graph structure provided with each dataset.
Appendix D Additional details on CoRel implementation
As discussed in Sec. 3 and Sec. 5, we implemented CoRel as TTS model with a single-layer GRU followed by $2$ message passing layers. As a readout, we used an MLP mapping the learned representations to prediction of $39$ equally spaced quantiles.
The graph learning module was implemented as described in Sec. 3.2 by parametrizing ${\bm{\Phi}}$ with a matrix of $N\times N$ learnable parameters. To allow for sampling less neighbors than the specified neighborhood size $K$, we followed previous works [40] and modified the sampling procedure by introducing a set of dummy nodes then discarded from the sampled graph before message passing. We use dummy nodes in the GPVAR experiment only. To allow for sparse message-passing operations we use a straight-through gradient estimator [45] and backpropagate gradient only through the sampled edges for each node plus $10\%$ of the remaining ones. In the GPVAR experiment we simply propagate gradients w.r.t. the entire adjacency matrix.
Besides residuals, we use as additional covariates datetime encodings whenever available plus the value of the target time series w.r.t. the time steps in the input window.
Appendix E Evaluation metrics
For a prediction interval $\widehat{C}_{i}^{\alpha}=[\hat{x}_{i}+\hat{q}_{i}^{\alpha/2},\hat{x}_{i}+\hat{q}_{i}^{1-\alpha/2}]$ and true value $x_{i}$, evaluation is conducted for a desired confidence level $\alpha$ using three key metrics.
Coverage gap
$\Delta\text{Cov}_{i}$ measures the difference between the achieved coverage and the target coverage $1-\alpha$, and is given by
[…]
where $\mathbbm{1}$ denotes the indicator function.
Prediction interval width
Quantifies the width of the prediction intervals, and is given by
[…]
Winkler Score
Combines interval width with a penalty for predictions that fail to capture the true value, with misses penalized by a factor of $\frac{2}{\alpha}$. The Winkler score is given by
$W_{i}=\begin{cases}(\hat{q}^{1-\alpha/2}_{i}-\hat{q}^{\alpha/2}_{i})+\frac{2}{\alpha}(\hat{q}^{\alpha/2}_{i}-x_{i})&\text{if }x_{i}<\hat{q}^{\alpha/2}_{i},\\ (\hat{q}^{1-\alpha/2}_{i}-\hat{q}^{\alpha/2}_{i})&\text{if }\hat{q}^{\alpha/2}_{i}\leq x_{i}\leq\hat{q}^{1-\alpha/2}_{i},\\ (\hat{q}^{1-\alpha/2}_{i}-\hat{q}^{\alpha/2}_{i})+\frac{2}{\alpha}(x_{i}-\hat{q}^{1-\alpha/2}_{i})&\text{if }x_{i}>\hat{q}^{1-\alpha/2}_{i}.\end{cases}$ (26)
All the metrics are then averaged over all nodes and time steps within the specified set.
Appendix F Hyperparameters and empirical setup
Hyperparameters were tuned separately for each combination of base predictor and dataset on a validation set.
CoRel
For CoRel we tuned the number of neurons in the STGNN with a small grid search on $10\%$ of the calibration data. For the experiments on real-word data, the model was training for a maximum of $100$ epochs the calibration set. Each epoch consisted of a maximum of $50$ mini-batches of size $64$. We used the Adam optimizer [80] with an initial learning rate of $0.003$ and reduced by $75\%$ every $20$ epochs. We used a fixed number of $K=20$ neighbors for the grapn learning module. We used the same hyperparameters and model selection procedure for CoRNN.
HopCPT
For HopCPT we followed the model selection procudere described in [21]. The model was trained for $3000$ epochs using a batch size of $4$ time series. We adopted the paper’s AdamW optimizer configuration with standard parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=0.01$) and tuned the model by running the same hyperparameter configurations searched by [21]. All the remaining hyperparameters were set accordingly to the original paper.
SeqCP
[…]
NexCP
[…]
SCPI
As already mentioned, we used the SCPI [19] implementation provided by Auer et al. [21] and followed an analogous protocol for training the mode. In particular, SCPI was run using a fixed window length of $100$ time steps for all experiments, corresponding to the longest setting in the original paper. The computational demands of SPCI were substantial, as separate quantile random forests had to be trained for each combination of node and target coverage level, making extensive hyperparameter tuning impractical. To further manage the computational costs, we trained each SPCI model only once on the calibration data, rather than implementing the time-adaptive approach where models are re-trained as new observations become available.
Appendix G Optimization of the PI width
Here, we show how the width of the PI can be reduced by searching for an appropriate quantile offset $\beta$. Specifically, we perform the optimization described in Eq. 16, which identifies pairs of quantiles yielding the same coverage $1-\alpha$ while achieving the smallest interval width. We summarize our results in Tab. 4. As we can see, the procedure consistently finds intervals with a smaller width while retaining approximately the same coverage, whether using CoRel or the CoRNN variant.
Table 4: Changes in $\Delta$ Cov and PI-Width when optimizing $\beta$ ($\alpha=0.1$).
RNN Transformer STGNN
$\Delta$Cov PI-Width $\Delta$Cov PI-Width $\Delta$Cov PI-Width
METR-LA CoRNN -1.52 $\pm$ 0.35 19.26 $\pm$ 0.27 -1.01 $\pm$ 0.28 19.45 $\pm$ 0.13 -0.40 $\pm$ 0.57 16.23 $\pm$ 0.36
$\beta$ optim. -1.50 $\pm$ 0.32 17.17 $\pm$ 0.21 -1.28 $\pm$ 0.18 17.16 $\pm$ 0.17 -0.64 $\pm$ 0.44 14.95 $\pm$ 0.32
CoRel -1.15 $\pm$ 0.48 15.67 $\pm$ 0.19 -0.47 $\pm$ 0.33 16.14 $\pm$ 0.38 -0.80 $\pm$ 0.49 15.73 $\pm$ 0.37
$\beta$ optim. -1.40 $\pm$ 0.52 14.58 $\pm$ 0.15 -0.91 $\pm$ 0.43 15.00 $\pm$ 0.33 -0.89 $\pm$ 0.48 14.82 $\pm$ 0.39
CER-E CoRNN -2.12 $\pm$ 0.19 2.27 $\pm$ 0.01 -2.05 $\pm$ 0.12 2.25 $\pm$ 0.01 -2.19 $\pm$ 0.13 2.04 $\pm$ 0.02
$\beta$ optim. -2.41 $\pm$ 0.15 2.20 $\pm$ 0.01 -2.31 $\pm$ 0.08 2.19 $\pm$ 0.01 -2.47 $\pm$ 0.13 2.00 $\pm$ 0.02
CoRel -3.56 $\pm$ 0.24 1.92 $\pm$ 0.01 -3.10 $\pm$ 0.33 1.95 $\pm$ 0.01 -4.90 $\pm$ 0.34 1.82 $\pm$ 0.03
$\beta$ optim. -4.09 $\pm$ 0.27 1.85 $\pm$ 0.01 -3.38 $\pm$ 0.32 1.87 $\pm$ 0.01 -5.28 $\pm$ 0.36 1.75 $\pm$ 0.03
AQI CoRNN -0.42 $\pm$ 0.84 75.85 $\pm$ 2.13 -0.10 $\pm$ 0.46 77.37 $\pm$ 1.32 -0.17 $\pm$ 0.42 $7$$\pm$ 0.43
$\beta$ optim. -0.36 $\pm$ 0.85 72.96 $\pm$ 1.96 0.05 $\pm$ 0.42 74.73 $\pm$ 1.16 -0.12 $\pm$ 0.40 75.58 $\pm$ 0.43
CoRel -1.53 $\pm$ 1.60 70.47 $\pm$ 3.60 -2.93 $\pm$ 1.76 71.04 $\pm$ 4.76 -2.50 $\pm$ 1.12 71.88 $\pm$ 2.87
$\beta$ optim. -2.78 $\pm$ 1.79 67.30 $\pm$ 3.37 -3.70 $\pm$ 1.79 67.95 $\pm$ 4.31 -3.51 $\pm$ 1.04 69.13 $\pm$ 2.73